# Ordinary Least Squares as an Attention Mechanism

### v1

@@ 1 Introduction
The self-attention mechanism popularized by vaswani2017attention has sparked a technical revolution that fundamentally reshaped the landscape of deep learning and propelled AI into the public eye. Its introduction eight years ago single-handedly ushered in the era of large language models (LLMs), shifting the natural language processing field away from convolutional/recurrent neural network architectures toward Transformer-based models.
@@
Outline. The rest of this paper is organized as follows. Section 2 presents an alternative view on OLS, reviews the attention mechanism, and connects the two under some conditions. Section 3 discusses extensions of these ideas to incorporate many other features of the actual attention mechanism employed in LLMs. Section 4 leverages the nonlinear regression interpretation of attention to unpack key elements of the Transformer architecture beyond the attention module itself, drawing connections to time series econometrics. Section 5 concludes by discussing some of the implications and suggesting directions for future research.
@@ 3.2 Nonlinearity and the Use of Softmax
In light of this, the Attention Regression emerges as a possibly useful hybrid, combining smooth nonlinearities found in kernel methods and neural networks with the simplex constraint on weights, a feature more commonly seen in tree-based models. Whether this translates into meaningful gains in predictive accuracy is an application-dependent empirical question.
@@ 5 Discussion and Directions for Future Research
The “OLS perspective” suggests that attention mechanisms can be understood as a nonlinear least squares method with generated features, where learned concepts replace raw predictors. These concepts emerge from a dimension-reduced representation of the input, akin to transformed tokens or feature embeddings. Instead of forming direct mappings, the attention mechanism restructures prediction as a correlation-based problem in this learned space, much like any nonparametric supervised learning approach. [^2] This view naturally aligns with the rest of the deep learning canon, where learned latent or hidden representations are a hallmark of the model’s ability to capture complex patterns. In certain contexts, these latent concepts can be extracted and visualized to enhance interpretability (kim2018interpretability; goulet2024neural).
The output of an attention module, often termed a contextualized embedding, integrates information across a sequence through projection rather than explicit sequential modeling. This aligns with time series econometrics tools like VARs, where equations can be estimated separately via OLS as if they were cross-sectional data, once lags are encoded. Each token’s contextualized embedding can be seen as a fitted value in a model where tokens simultaneously “explain” and are “explained by” others.
A macroeconomic parallel is the interpretation of high interest rates across inflationary environments: their meaning shifts depending on context, much like attention-based embeddings adjust dynamically. Context for interest rates is often formalized through a Taylor rule, where higher inflation and stronger real activity justify higher rates. Rates that align with the rule carry different implications than those that deviate significantly—known as monetary policy shocks. Again, “context” takes the form of expected conditions at $t$ given available information.
@@
Macroeconomists are particularly fond of shocks, the difference between the realized value of $y_{t}$ and its expected value. In contrast, LLMs construct conditional means as indicators of context, serving as a form of feature engineering for subsequent feed-forward neural network modules. In this spirit, it is intriguing to consider the possibility of word shocks. One could examine the difference between the original, context-independent embeddings and their context-conditioned versions. This difference could be interpreted as a “shock” or a granular measure of linguistic surprise–capturing unexpected word choices or unusual word adjacencies. Some factor model or supervised aggregation scheme (as in, e.g., albacore) could be leveraged to map this back into macroeconomic series of interest.
This granular approach would differ significantly from methods that construct a single shock series by regressing an observed time series (such as short-term interest rates) on sentiment extracted from text (aruoba2024identifying). Instead, each word in a sentence could be decomposed into the sum of its contextual contribution and a shock component. Intermediate aggregation schemes could also prove useful, such as ranking policy reports on a spectrum from “utterly predictable” to “jarring.” Some existing work moves in this direction, such as fischer2023fed, who tracks various aspects of Fed communications, including their opacity level.

### v2

@@ 1 Introduction
The self-attention mechanism popularized by vaswani2017attention has sparked a technical revolution that fundamentally reshaped the landscape of deep learning and propelled AI into the public eye. Its introduction in 2017 single-handedly ushered in the era of large language models (LLMs), shifting the natural language processing field away from convolutional/recurrent neural network architectures toward Transformer-based models.
@@
Outline. The rest of this paper is organized as follows. Section 2 presents an alternative view on OLS, reviews the attention mechanism, and connects the two under some conditions. Section 3 discusses extensions of these ideas to incorporate many other features of the actual attention mechanism employed in LLMs. Section 4 leverages the nonlinear regression interpretation of attention to unpack key elements of the Transformer architecture beyond the attention module itself, drawing connections to time series econometrics. Section 5 provides experimental evidence on the predictive performance of the nonlinear Attention Regression framework across a range of data-generating processes. Section 6 concludes by discussing some of the implications and suggesting directions for future research.
@@ 2.4 The OLS-Attention Nexus
Putting it all together.
Figure 1 summarizes the correspondence between OLS predictions and the attention mechanism. The test observations $\boldsymbol{X}_{\text{test}}$ and training observations $\boldsymbol{X}_{\text{train}}$ are projected into an orthonormal space via the shared encoding $\boldsymbol{W}=\boldsymbol{U}\Lambda^{-1/2}$, yielding queries $\mathbf{Q}$ and keys $\mathbf{K}$, respectively. The inner product $\mathbf{Q}\mathbf{K}^{\prime}$ produces a similarity matrix whose entries $\omega_{ji}$ weight the training outcomes $\boldsymbol{y}$ to form predictions. In the linear case, these weights are unbounded and may take negative values—properties that distinguish OLS from softmax-based attention, where weights are constrained to the simplex. The matrix $\boldsymbol{\Omega}=\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{\prime}=(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}})^{-1}$ emerges as the solution to an optimal embedding problem: it defines the latent geometry in which proximity-based weighting minimizes squared prediction error.
Figure 1: OLS Predictions as a Linear Attention Module
@@ 3.2 Nonlinearity and the Use of Softmax
In light of this, the Attention Regression emerges as a possibly useful hybrid, combining smooth nonlinearities found in kernel methods and neural networks with the simplex constraint on weights, a feature more commonly seen in tree-based models. Whether this translates into meaningful gains in predictive accuracy is an application-dependent empirical question. Section 5 takes up this question directly, evaluating the finite-sample performance of Attention Regression against standard nonlinear benchmarks across a range of data-generating processes.
@@ 4.5 Recursive Chaining of Attention Modules and Feed-Forward Networks
5 Experimental Evidence
We conduct a Monte Carlo experiment to evaluate the finite-sample properties of the attention-based regression framework discussed in Section 3.2. While I have established that OLS represents a specific case of linear attention, this experiment assesses the performance of the multi-head nonlinear attention regression across a range of functional forms and signal-to-noise environments.
Attention-Based Regression Specification.
I implement the multi-head extension of the nonlinear attention regression framework introduced in Section 3.2. For a matrix of test points $\boldsymbol{X}^{*}$ and training data $\boldsymbol{X}_{\text{train}}$ with outcomes $\boldsymbol{y}$, the estimator forms predictions by aggregating $M$ distinct attention mechanisms:
$\hat{\boldsymbol{y}}=\sum_{m=1}^{M}\alpha_{m}\left[\text{softmax}\left(\boldsymbol{X}^{*}\boldsymbol{\Omega}_{m}\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\right],$ (44)
where $\alpha_{m}$ are scalar, head-specific combination weights. Each head learns a unique similarity metric $\boldsymbol{\Omega}_{m}$, which I parameterize via the Cholesky factorization $\boldsymbol{\Omega}_{m}=\boldsymbol{L}_{m}\boldsymbol{L}_{m}^{\prime}$ (with $\boldsymbol{L}_{m}$ being lower triangular). This ensures that $\boldsymbol{\Omega}_{m}$ remains positive semi-definite, effectively mapping the data into a latent space where similarities are computed as valid inner products.
The model parameters $\{\alpha_{m},\boldsymbol{L}_{m}\}_{m=1}^{M}$ are jointly estimated by minimizing the penalized squared error loss:
$\mathcal{L}=\left\|\boldsymbol{y}-\sum_{m=1}^{M}\alpha_{m}\text{softmax}\left(\boldsymbol{X}_{\text{train}}(\boldsymbol{L}_{m}\boldsymbol{L}_{m}^{\prime})\boldsymbol{X}_{\text{train}}^{\prime}\right)\boldsymbol{y}\right\|^{2}+\lambda\sum_{m=1}^{M}\|\boldsymbol{L}_{m}\|_{F}^{2}.$ (45)
Optimization is performed via L-BFGS with early stopping. To anchor the nonconvex optimization near a sensible starting point, each $\boldsymbol{L}_{m}$ initialized using the Cholesky factor of the ridge-regularized precision matrix $(\boldsymbol{X}_{\text{train}}^{\prime}\boldsymbol{X}_{\text{train}}+\lambda\boldsymbol{I})^{-1}$, scaled by $\sqrt{N}$, with added Gaussian noise to promote diversity across heads.
Data-Generating Processes.
Let $\boldsymbol{X}_{\text{train}}\in\mathbb{R}^{N\times 5}$ with rows drawn independently from $\text{Uniform}[0,1]^{5}$. For each DGP, outcomes are generated according to
$y_{i}=f(\boldsymbol{X}_{i})+\varepsilon_{i},\qquad\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),$
where $\sigma^{2}=\text{Var}(f(\boldsymbol{X}_{i}))/\text{SNR}$ calibrates the noise to a target signal-to-noise ratio. I employ six standard regression functions to test the model’s adaptability:
  1. 1.
    Linear: $f(\boldsymbol{X}_{i})=2(X_{i,1}-0.5)-(X_{i,2}-0.5)+3(X_{i,3}-0.5)+1.5(X_{i,4}-0.5)+0.5(X_{i,5}-0.5).$
  2. 2.
    Friedman 1: $f(\boldsymbol{X}_{i})=10\sin(\pi X_{i,1}X_{i,2})+20(X_{i,3}-0.5)^{2}+10X_{i,4}+5X_{i,5}.$
  3. 3.
    Friedman 2: $f(\boldsymbol{X}_{i})=\sin\bigl(\pi(X_{i,1}+X_{i,2}+X_{i,3})\bigr)+\log(1+X_{i,4}^{2}).$
  4. 4.
    Friedman 3: $f(\boldsymbol{X}_{i})=X_{i,1}X_{i,2}+\log(X_{i,3}+X_{i,4}+2).$
  5. 5.
    Rotated Sine: $f(\boldsymbol{X}_{i})=\sin\Bigl(3\sum_{j=1}^{4}X_{i,j}\Bigr).$
  6. 6.
    Soft Radial: $f(\boldsymbol{X}_{i})=\bigl(1+5\|\boldsymbol{X}_{i}-0.5\boldsymbol{1}\|^{2}\bigr)^{-1}.$
These functions span linear, additive nonlinear, interactive, periodic, and radial structures, providing a diverse testbed that does not privilege any specific estimator.
Design and Competitors.
The training sample size is $N\in\{500,1000,2500,5000\}$ and signal-to-noise ratio $\text{SNR}\in\{0.5,1,2,3\}$. A fixed test set of $J=1000$ observations is used throughout, with results averaged over 10 replications. I compare five methods: ordinary least squares (OLS), Random Forest with 500 trees and $\lfloor P/3\rfloor$ variables per split, a Multi-Layer Perceptron (MLP) with three hidden layers of 200 units each, ReLU activations, dropout regularization (rate =0.2=0.2), trained via Adam with early stopping (patience =20=20 epochs, 15% validation holdout), Gradient Boosting with 500 trees and learning rate 0.01, and Attention Regression with five heads and regularization $\lambda=10^{-3}$. Performance is measured by out-of-sample $R^{2}$, which is informative because it is comparable across DGPs and can be analyzed in relation to the maximal theoretically attainable $R^{2}$, which is $R^{2}_{\max}=\text{SNR}/(\text{SNR}+1)$.
Figure 2: Model Performance Comparison Across Data Generating Processes
Notes: This figure displays the average $R^{2}$ values for each model across all experimental conditions (sample sizes $N\in\{500,1000,2500,5000\}$ and signal-to-noise ratios $\text{SNR}\in\{0.5,1.0,2.0,3.0\}$). Error bars represent $\pm 1$ standard deviation across the 16 conditions within each DGP. The “Overall” category aggregates performance across all six DGPs (96 total conditions). Values above bars indicate the mean $R^{2}$ of the best-performing model for each DGP. The dashed vertical line separates the aggregate results from individual DGP performance.
Results.
As summarized in Figure 2 and detailed in Table LABEL:tab:benchmark_results, the MLP achieves the highest overall $R^{2}$ (0.52) across all 96 experimental conditions, followed by Attention Regression (0.48), Random Forest (0.45), Gradient Boosting (0.42), and OLS (0.28). The MLP’s strong performance is expected: as a universal function approximator with flexible hidden representations, it can adapt to arbitrary nonlinearities given sufficient data.
What is more notable is how competitive Attention Regression remains despite its simpler architecture. It consistently outperforms both Random Forest and Gradient Boosting overall and in four of the six DGPs. On Friedman1, it wins outright (0.53 vs. 0.52 for MLP and 0.49 for RF/GBM), and it essentially ties the MLP on Friedman3 (0.54 for both) while beating RF and GBM by 4 and 1 percentage points respectively. On Friedman2, Attention Regression outperforms Random Forest by 7 percentage points and Gradient Boosting by 15 points, though the MLP leads by about 5 points. These Friedman designs rely heavily on interactions and smooth nonlinear transformations—exactly the patterns that learned attention weights can capture through adaptive similarity weighting.
The expected exception is the Linear DGP, where OLS dominates (0.56) due to correct specification. Here, all flexible methods pay a price for their flexibility, with Attention Regression trailing by about 8 points. On the periodic and radial DGPs (Rotated Sine and Soft Radial), Attention Regression beats both tree-based methods on Rotated Sine (0.41 vs. 0.38 for RF and 0.25 for GBM) while the MLP’s advantage is more pronounced, reflecting its ability to approximate high-frequency oscillations through compositions of ReLU activations. On Soft Radial, Random Forest leads among non-neural methods (0.47), with Attention Regression (0.46) close behind.
As visible from the error bars in Figure 2, variability is higher for Attention Regression and MLP in small samples or low SNR settings, but this gap narrows with larger $N$, consistent with typical sample-efficiency patterns of flexible nonlinear models.
These simulations demonstrate that a straightforward implementation of Attention Regression performs remarkably well relative to standard nonlinear ML benchmarks. All models were run with basic, off-the-shelf hyperparameters—the purpose of the paper is not to introduce yet another nonlinear estimator, but to clarify the connection between attention mechanisms and least squares. Still, it is notable that taking the formulation literally and estimating $\boldsymbol{\Omega}$ directly produces a model that often matches tree-based methods and remains competitive with neural networks.
@@ 6 Discussion and Directions for Future Research
The "OLS perspective" suggests that attention mechanisms can be understood as a nonlinear least squares method with generated features, where learned concepts replace raw predictors. These concepts emerge from a dimension-reduced representation of the input, akin to transformed tokens or feature embeddings. Instead of forming direct mappings, the attention mechanism restructures prediction as a correlation-based problem in this learned space, much like any nonparametric supervised learning approach. [^2] This view naturally aligns with the rest of the deep learning canon, where learned latent or hidden representations are a hallmark of the model’s ability to capture complex patterns. In certain contexts, these latent concepts can be extracted and visualized to enhance interpretability (kim2018interpretability; goulet2024neural).
The output of an attention module, often termed a contextualized embedding, integrates information across a sequence through projection rather than explicit sequential modeling. This aligns with time series econometrics tools like VARs, where equations can be estimated separately via OLS as if they were cross-sectional data, once lags are encoded. Each token’s contextualized embedding can be seen as a fitted value in a model where tokens simultaneously "explain" and are "explained by" others.
A macroeconomic parallel is the interpretation of high interest rates across inflationary environments: their meaning shifts depending on context, much like attention-based embeddings adjust dynamically. Context for interest rates is often formalized through a Taylor rule, where higher inflation and stronger real activity justify higher rates. Rates that align with the rule carry different implications than those that deviate significantly—known as monetary policy shocks. Again, "context" takes the form of expected conditions at $t$ given available information.
@@
Macroeconomists are particularly fond of shocks, the difference between the realized value of $y_{t}$ and its expected value. In contrast, LLMs construct conditional means as indicators of context, serving as a form of feature engineering for subsequent feed-forward neural network modules. In this spirit, it is intriguing to consider the possibility of word shocks. One could examine the difference between the original, context-independent embeddings and their context-conditioned versions. This difference could be interpreted as a "shock" or a granular measure of linguistic surprise–capturing unexpected word choices or unusual word adjacencies. Some factor model or supervised aggregation scheme (as in, e.g., albacore) could be leveraged to map this back into macroeconomic series of interest.
This granular approach would differ significantly from methods that construct a single shock series by regressing an observed time series (such as short-term interest rates) on sentiment extracted from text (aruoba2024identifying). Instead, each word in a sentence could be decomposed into the sum of its contextual contribution and a shock component. Intermediate aggregation schemes could also prove useful, such as ranking policy reports on a spectrum from "utterly predictable" to "jarring." Some existing work moves in this direction, such as fischer2023fed, who tracks various aspects of Fed communications, including their opacity level.
@@
Recent Developments.
Since the initial circulation of this paper, the landscape of large language models has continued to evolve rapidly. Reasoning-focused architectures and retrieval-augmented generation (RAG) systems have gained prominence, yet the attention mechanism remains their computational backbone. Alternative architectures such as Mamba (gu2023mamba) and FFT-based methods (fein2025fft) have emerged as potential successors, but attention-based Transformers continue to dominate deployed systems. The OLS perspective developed here remains applicable to these newer variants insofar as they rely on similarity-based weighting of stored representations or can be framed as adaptive filters in a latent space.
@@
Appendix A Detailed Simulation Results
Table 1: Model Performance Comparison (Out-of-Sample $R^{2}$) Across Data Generating Processes
Models
DGP NN SNR OLS RF MLP GBM Att Reg
[…]
1000 0.5 0.32 0.27 0.31 0.28 0.29
1 0.49 0.44 0.47 0.45 0.42
2 0.66 0.60 0.64 0.62 0.55
3 0.74 0.68 0.73 0.71 0.68
[…]
5000 0.5 0.34 0.31 0.33 0.32 0.26
1 0.50 0.47 0.49 0.48 0.38
2 0.67 0.64 0.66 0.65 0.52
3 0.75 0.72 0.75 0.73 0.57
[…]
1000 0.5 0.25 0.27 0.29 0.27 0.30
1 0.37 0.42 0.45 0.43 0.47
2 0.50 0.58 0.62 0.59 0.64
3 0.56 0.66 0.71 0.67 0.73
[…]
5000 0.5 0.25 0.30 0.32 0.30 0.33
1 0.37 0.46 0.48 0.45 0.48
2 0.49 0.62 0.65 0.61 0.64
3 0.55 0.70 0.73 0.68 0.72
[…]
1000 0.5 0.02 0.18 0.27 0.12 0.23
1 0.04 0.30 0.44 0.23 0.41
2 0.05 0.42 0.61 0.34 0.57
3 0.06 0.49 0.70 0.40 0.63
[…]
5000 0.5 0.04 0.26 0.31 0.20 0.27
1 0.05 0.41 0.47 0.30 0.41
2 0.07 0.55 0.64 0.41 0.52
3 0.08 0.63 0.73 0.46 0.63
[…]
1000 0.5 0.30 0.27 0.30 0.29 0.31
1 0.45 0.43 0.47 0.46 0.48
2 0.60 0.59 0.64 0.63 0.65
3 0.67 0.67 0.73 0.71 0.74
[…]
5000 0.5 0.29 0.29 0.31 0.31 0.31
1 0.44 0.46 0.48 0.47 0.48
2 0.59 0.63 0.65 0.64 0.65
3 0.66 0.71 0.73 0.72 0.74
[…]
1000 0.5 0.10 0.18 0.27 0.12 0.21
1 0.15 0.30 0.43 0.20 0.37
2 0.20 0.42 0.61 0.30 0.51
3 0.23 0.48 0.70 0.35 0.56
[…]
5000 0.5 0.10 0.26 0.30 0.16 0.22
1 0.16 0.40 0.47 0.26 0.36
2 0.21 0.54 0.64 0.35 0.47
3 0.24 0.61 0.73 0.40 0.58
[…]
1000 0.5 -0.00 0.26 0.27 0.23 0.21
1 -0.00 0.40 0.45 0.36 0.38
2 -0.00 0.55 0.62 0.49 0.55
3 -0.00 0.62 0.70 0.56 0.63
[…]
5000 0.5 -0.00 0.29 0.31 0.25 0.29
1 -0.00 0.45 0.48 0.39 0.45
2 -0.00 0.61 0.65 0.52 0.56
3 -0.00 0.69 0.73 0.59 0.62
Notes: Each cell reports out-of-sample $R^{2}$. Models: OLS (Ordinary Least Squares), RF (Random Forest), MLP (Multi-Layer Perceptron), GBM (Gradient Boosting), Att Reg (Attention Regression). Green bold = best; black bold = second-best.